# RBI-ThPhys-2024-16 ZTF-EP-24-07

### v2

@@ I Introduction
At this energy scale, the gravitational effects become comparable to quantum effects and one might expect a modification of GR to be triggered there. This is the point where the gravitational waves come into play. Besides representing one of the most important predictions of general relativity, gravitational waves appear to play a crucial role in accessing the behaviour of GR in extreme conditions under which it is yet not so well studied. Namely, due to the fact that gravitational waves originate from regions of spacetime where gravity is extremely strong, just like the near-horizon region of black holes, they came out as the most suitable candidate for testing GR under such extreme conditions. In fact, since their recent experimental discovery, gravitational waves have been established as one of the most powerful tools for studying extreme regimes of relativistic speeds, large curvature, and strong gravity. In this way, gravitational waves provided a promising way for exploring spacetime dynamics in regions previously inaccessible to an experimental enquiry and helped bring them under a direct experimentalist’s gaze.
@@
Another related observation involves GR itself and its validity and perseverance in the vicinity of a black hole. As already stated before, the region near a black hole horizon admits access to energies of the order of the Planck scale. As a result, GR may be potentially modified there, causing that some of its predictions, including gravitational quasinormal mode (QNM) spectrum of black holes, be altered too [5, 6, 7, 8, 9]. The latter scenario receives significant support within a scope of various quantum gravity theories, which suggest that strong gravity and high energy regimes like those found in a near-horizon region of black hole may create environment where the usual point-like structure of spacetime becomes replaced with certain noncommutative geometry constructs [10, 11] ultimately affecting the relaxation dynamics of black holes and their related ringdown spectra. Even more striking evidence of the discrete nature of spacetime and a possible necessity for modifying classical GR might be found in phenomena like gravitational wave echoes or tidal heating [12].
@@ II.2 Quantum Lie algebras and NC geometry
$\hat{R}(v,w,z):=$$\>\hat{\nabla}_{v}\hat{\nabla}_{w}z-\hat{\nabla}_{\bar{R}^{\alpha}(w)}\hat{\nabla}_{\bar{R}^{\alpha}(v)}z-\hat{\nabla}_{[v,w]_{\star}}z,$
@@
$\hat{R}(\partial_{\mu},\partial_{\nu},\partial_{\rho})=$ $\ \hat{R}_{~{}\mu\nu\rho}^{\sigma}=\left(\partial_{\mu}\hat{\Gamma}_{~{}\nu\rho}^{\sigma}-\partial_{\nu}\hat{\Gamma}_{~{}\mu\rho}^{\sigma}+\hat{\Gamma}_{~{}\nu\rho}^{\tau}\star\hat{\Gamma}_{~{}\mu\tau}^{\sigma}-\hat{\Gamma}_{~{}\mu\rho}^{\tau}\star\hat{\Gamma}_{~{}\nu\tau}^{\sigma}\right)\partial_{\sigma},$
@@ III.1 Axial perturbations and noncommutative Regge-Wheeler potential
The radial parts of the linearized equations $\hat{{\rm R}}_{r\theta}=0$, $\hat{{\rm R}}_{\theta\theta}=0$, $\hat{{\rm R}}_{\varphi\varphi}=0$, and $\hat{{\rm R}}_{t\theta}=0$ are identical to the above equations. Among the 3 equations (34-36), only 2 are mutually independent. We are therefore in position to pick any two out of these three equations to find the solution of the whole system, the easiest choice being first-order equations (34) and (36). Combining these equations eliminates $h_{0}$ and yields a single second-order differential equation for $h_{1}$, as demonstrated in [19],
@@
Finally, we arrive at the differential equation,
@@ III.2 Polar perturbations and noncommutative Zerilli potential
where we used the abbreviation $\ell(\ell+1)=2\Lambda+2$. Note that the noncommutative part of the tortoise coordinate differs from that of axial perturbation (see Ref. [19]) since it depends on both $\ell$ and $m$. The resulting Schrödinger-like form of the equation (59) features an effective potential expressed as $V=V_{\text{Z}}+\,V_{\text{NC}}$, where $V_{\text{Z}}$ denotes the commutative Zerilli potential, and $V_{\text{NC}}$ represents the noncommutative correction. The explicit forms of these potentials are
@@ VI Outlook and discussion
The analysis has been made in a bottom-up approach, implying that the equation of motion is obtained by carefully revising the basic notions of differential geometry, such as the vector field, Lie derivative, covariant derivative, etc, rather than from some anticipated overruling/paramount action principle. In particular, a deformation quantization by a Drinfeld twist operator has been applied to a linearized gravitational perturbation theory. After developing a general perturbation theory up to linear order in both the deformation parameter and metric perturbation, we have applied it to a fixed Schwarzschild background in order to analyze the corresponding axial as well as polar type of gravitational perturbations. In this way, we have derived the noncommutative generalizations of the Regge-Wheeler and Zerilli potentials, respectively.
@@ Appendix A Zerilli gauge
$h_{rr}=\frac{1}{A(r)}\sum_{\ell,m}H_{2}^{\ell m}(t,r)Y_{\ell m}(\theta,\varphi),$ (82)
$h_{ta}=\sum_{\ell,m}\beta^{\ell m}(t,r)\partial_{a}Y_{\ell m}(\theta,\varphi),\quad h_{ra}=\sum_{\ell,m}\alpha^{\ell m}(t,r)\partial_{a}Y_{\ell m}(\theta,\varphi),$ (83)
@@
In the polar sector, one can choose the gauge parameter $\xi$ in such a way that three families of functions in (81 - 84), namely $G^{\ell m}$, $\alpha^{\ell m}$, and $\beta^{\ell m}$, are set to zero. This particular gauge is known as the Zerilli gauge. In the Zerilli gauge, the polar perturbations can be parametrized by the remaining four families of functions, namely $H_{0}^{\ell m}$, $H_{1}^{\ell m}$, $H_{2}^{\ell m}$, and $K^{\ell m}$, as follows:
@@ Appendix C Error estimation in WKB approximation method
In this appendix, we provide the details of error estimation in the QNM calculation using the higher-order WKB method for both axial and polar modes. Error for the $n$-th order WKB is equal to $|\omega_{n+1}-\omega_{n-1}|/2$ [34]. For the commutative case $(am=0)$ with $\ell=2$ we obtain

### v3

@@ I Introduction
At this energy scale, the gravitational effects become comparable to quantum effects and one might expect a modification of GR to be triggered there. This is one of those points where the gravitational waves come into play. Besides them being one of the most important predictions of general relativity, gravitational waves appear to play a crucial role in accessing the behaviour of GR under extreme conditions, i.e. in a regime where GR is not well enough studied yet. Namely, due to the fact that gravitational waves originate from regions of spacetime where gravity is extremely strong, just like the near-horizon region of black holes, they have emerged as highly suitable candidates for testing GR under such extreme conditions. In fact, since their recent experimental discovery, gravitational waves have been established as one of the most powerful tools for studying extreme regimes of relativistic speeds, large curvature, and strong gravity. In this way, gravitational waves provided a promising way for exploring spacetime dynamics in regions previously inaccessible to an experimental enquiry and helped bring them under a direct experimentalist’s gaze.
@@
However, strong gravity regimes, such as those found in the vicinity of a black hole, may present a more challenging scenario for general relativity, as its validity could be subjected to a more rigorous examination in these contexts. As already stated, the region near a black hole horizon admits access to energies of the order of the Planck scale. As a result, GR may be potentially modified there, causing that some of its predictions, including gravitational QNM spectrum of black holes, may be altered too [5, 6, 7, 8, 9, 10]. Indeed, the latter scenario receives significant support within a scope of various quantum gravity theories, which suggest that strong gravity and high energy regimes like those found in a near-horizon region of black hole may create environment where the usual point-like structure of spacetime becomes replaced with certain noncommutative geometry constructs [11, 12] ultimately affecting the relaxation dynamics of black holes and their related ringdown spectra. Even more striking evidence of the discrete nature of spacetime and a possible necessity for modifying classical GR might be found in phenomena like gravitational wave echoes or tidal heating [13]. Alternative theories of gravity – such as the Einstein-Aether model [14], Einstein-power-Maxwell scenarios [15, 16] and studies of acoustic black holes [17] – have similarly revealed that deviations from the classical GR QNM spectra may occur.
@@ II.2 Quantum Lie algebras and NC geometry
$\hat{R}(v,w,z):=$$\>\hat{\nabla}_{v}\hat{\nabla}_{w}z-\hat{\nabla}_{\bar{R}^{\alpha}(w)}\hat{\nabla}_{\bar{R}_{\alpha}(v)}z-\hat{\nabla}_{[v,w]_{\star}}z,$
@@
$\hat{R}(\partial_{\mu},\partial_{\nu},\partial_{\rho})=$ $\ \hat{R}_{~{}\mu\nu\rho}^{\sigma}\partial_{\sigma}=\left(\partial_{\mu}\hat{\Gamma}_{~{}\nu\rho}^{\sigma}-\partial_{\nu}\hat{\Gamma}_{~{}\mu\rho}^{\sigma}+\hat{\Gamma}_{~{}\nu\rho}^{\tau}\star\hat{\Gamma}_{~{}\mu\tau}^{\sigma}-\hat{\Gamma}_{~{}\mu\rho}^{\tau}\star\hat{\Gamma}_{~{}\nu\tau}^{\sigma}\right)\partial_{\sigma},$
@@ III.1 Axial perturbations and noncommutative Regge-Wheeler potential
The radial parts of the linearized equations $\hat{{\rm R}}_{r\theta}=0$, $\hat{{\rm R}}_{\theta\theta}=0$, $\hat{{\rm R}}_{\varphi\varphi}=0$, and $\hat{{\rm R}}_{t\theta}=0$ are identical to the above equations. Among the 3 equations (34-36), only 2 are mutually independent. We are therefore in position to pick any two out of these three equations to find the solution of the whole system, the easiest choice being first-order equations (34) and (36). Combining these equations eliminates $h_{0}$ and yields a single second-order differential equation for $h_{1}$.
@@
Expressions for the tortoise coordinate and field redefinition come from two algebraic conditions that we have to impose in order to fix the Schrödinger form of the equation. Detailed derivation is given in Appendix C of [24]. Finally, we arrive at the differential equation,
@@ III.2 Polar perturbations and noncommutative Zerilli potential
where we used the abbreviation $\ell(\ell+1)=2\Lambda+2$. Similar to the axial case, the tortoise coordinate and field redefinition come from the two conditions (58) that we impose in order to fix the Schrödinger form of the equation. Note that the noncommutative part of the tortoise coordinate differs from that of axial perturbation (see Ref. [24]) since it depends on both $\ell$ and $m$. The resulting Schrödinger-like form of the equation (59) features an effective potential expressed as $V=V_{\text{Z}}+\,V_{\text{NC}}$, where $V_{\text{Z}}$ denotes the commutative Zerilli potential, and $V_{\text{NC}}$ represents the noncommutative correction. The explicit forms of these potentials are
@@ V Axial vs polar perturbations – violation of classical isospectrality
Polar modes exhibit both higher and lower imaginary QNM frequencies compared to axial modes, as seen in Figure 6. Specifically, for negative values of $am$, the polar modes have larger imaginary frequencies (implying stronger damping), while for positive values of $am$, they have smaller imaginary frequencies (weaker damping). This could lead to a more striking statistical data output on the polar modes side recorded at the gravitational wave detector, provided that the detector can distinguish between the polar and axial modes. Similarly, if the detector is able to distinguish azimuthal numbers of the waves and measure that $m$ and $-m$ have different damping times, this could hint at $r-\varphi$ quantization.
@@
The physical significance of breaking isospectrality is profound. In the context of gravitational wave observations, differing QNM frequencies can serve as a tool to distinguish between different gravitational theories. For instance, if future gravitational wave detections from black hole mergers reveal discrepancies in the expected QNM frequencies between axial and polar modes, it could indicate the presence of new physics beyond general relativity. Such observations may provide empirical tests of modified theories of gravity and offer insights into the fundamental nature of spacetime. Consequently, experimental bounds on isospectrality are directly tied to a possibility of deviating from general relativity. Detecting any such deviations in gravitational wave observations would provide critical information about the underlying quantum gravity effects, thus paving the way for new theoretical developments and a deeper understanding of the universe’s structure.
@@ VI Outlook and discussion
The analysis has been made in a bottom-up approach, implying that the equation of motion is obtained by carefully revising the basic notions of differential geometry, such as the vector field, Lie derivative, covariant derivative, etc, rather than from some anticipated overruling/paramount action principle. In particular, a deformation quantization by a Drinfeld twist operator has been applied to a linearized gravitational perturbation theory. After developing a general perturbation theory up to linear order in both the deformation parameter and metric perturbation, we have applied it to a fixed Schwarzschild background in order to analyze the corresponding axial as well as polar type of gravitational perturbations. In this way, we have derived the noncommutative generalizations of the Regge-Wheeler and Zerilli potentials, respectively. It should be noted that equations leading up to effective potential are valid for two sources of noncommutativity (and their linear combinations): $r-\varphi$ and $r-t$. The analysis of the spectrum has been carried out only for the $r-\varphi$ case.
The Abelian Drinfeld twist results in noncommutativity of Moyal type, which features simplicity, generality, and versatility at low-energy regimes, where it reflects general features of NC spacetime. Specifically, the semi-Killing form of the twist introduced non-trivial NC corrections in the perturbations around Schwarzschild spacetime while ensuring separability of equations. Alternative coordinate choices, involving non-constant functions, can still be expressed in a Moyal form when suitably adapted, thus emphasizing the versatility of Moyal twist-based approaches. Ultimately, our selected twist demonstrates explicitly how noncommutativity could modify gravitational wave signals, providing a specific, testable prediction and an important starting point for future experimental and theoretical inquiries into NC gravitational physics.
@@
A deeper insight into the spectrum and nature of the noncommutative metric perturbations can be gained by employing more robust methods, such as Leaver’s continued fraction method or direct integration, and by considering the time profile of the perturbations. It has been recently shown that even small perturbations to the spectrum can lead to instability [75]. We plan to address this in a future work. Preliminary steps have already been taken in [76], where we showed that another radial coordinate can be used to make the behaviour of the tortoise coordinate and effective potential more similar to commutative case in the near-horizon region, which is a necessary starting point for these more advanced methods.
@@ Appendix A Zerilli gauge
$h_{rr}=\frac{1}{A(r)}\sum_{\ell,m}H_{2}^{\ell m}(t,r)Y_{\ell m}(\theta,\varphi),$
$h_{ta}=\sum_{\ell,m}\beta^{\ell m}(t,r)\partial_{a}Y_{\ell m}(\theta,\varphi),\quad h_{ra}=\sum_{\ell,m}\alpha^{\ell m}(t,r)\partial_{a}Y_{\ell m}(\theta,\varphi),$
@@
In the polar sector, one can choose the gauge parameter $\xi$ in such a way that three families of functions in (83 - 84), namely $G^{\ell m}$, $\alpha^{\ell m}$, and $\beta^{\ell m}$, are set to zero. This particular gauge is known as the Zerilli gauge. In the Zerilli gauge, the polar perturbations can be parametrized by the remaining four families of functions, namely $H_{0}^{\ell m}$, $H_{1}^{\ell m}$, $H_{2}^{\ell m}$, and $K^{\ell m}$, as follows:
@@ Appendix C Error estimation in WKB approximation method
In this appendix, we provide the details of error estimation in the QNM calculation using the higher-order WKB method for both axial and polar modes. Error for the $n$-th order WKB is equal to $|\omega_{n+1}-\omega_{n-1}|/2$ [39, 80]. For the commutative case $(am=0)$ with $\ell=2$ we obtain